OpenStudy (missmeow):

A closed container has 4.02 ⋅ 10^23 atoms of a gas. Each atom of the gas weighs 1.67 ⋅ 10^−24 grams. Which of the following shows and explains the approximate total mass, in grams, of all the atoms of the gas in the container? 1) 6.71 grams, because (4.02 ⋅ 1.67) ⋅ (10^23 ⋅ 10^−24) = 6.7134 2) 5.69 grams, because (4.02 + 1.67) ⋅ (10^23 ⋅ 10^−24) = 5.69 3) 0.67 grams, because (4.02 ⋅ 1.67) ⋅ (10^23 ⋅ 10^−24) = 6.7134 ⋅ 10^−1 4) 0.57 grams, because (4.02 + 1.67) ⋅ (10^23 ⋅ 10^−24) = 5.69 ⋅ 10^−1 could you please explain how you got there? I want to learn!

OpenStudy (retireed):

First if each atoms is X, and each X weighs Y grams the total weight would be Atom 1 2 3 4 5 ..... X Total Number Atoms Grams Y+Y+Y+Y+Y ...... +Y+Y+Y Total Mass Instead of adding all of Atoms weight, it would be easier to multiply X times Y. Knowing this which two equations can be eliminated? 1) 6.71 grams, because (4.02 ⋅ 1.67) ⋅ (10^23 ⋅ 10^−24) = 6.7134 2) 5.69 grams, because (4.02 + 1.67) ⋅ (10^23 ⋅ 10^−24) = 5.69 3) 0.67 grams, because (4.02 ⋅ 1.67) ⋅ (10^23 ⋅ 10^−24) = 6.7134 ⋅ 10^−1 4) 0.57 grams, because (4.02 + 1.67) ⋅ (10^23 ⋅ 10^−24) = 5.69 ⋅ 10^−1

OpenStudy (missmeow):

I'm not sure

OpenStudy (missmeow):

2 and 4?

OpenStudy (retireed):

Which answers is using adding atoms to weight and which is multiplying atoms and weight? 4.02 ⋅ 10^23 atoms atom of the gas weighs 1.67 ⋅ 10^−24 grams 1) 6.71 grams, because (4.02 ⋅ 1.67) ⋅ (10^23 ⋅ 10^−24) = 6.7134 2) 5.69 grams, because (4.02 + 1.67) ⋅ (10^23 ⋅ 10^−24) = 5.69 3) 0.67 grams, because (4.02 ⋅ 1.67) ⋅ (10^23 ⋅ 10^−24) = 6.7134 ⋅ 10^−1 4) 0.57 grams, because (4.02 + 1.67) ⋅ (10^23 ⋅ 10^−24) = 5.69 ⋅ 10^−1

OpenStudy (missmeow):

1 and 3 are multiplying, and 2 and 4 are adding

OpenStudy (retireed):

Yes 2 and 4. I didn't see your second reply

OpenStudy (missmeow):

okay, what do I do next?

OpenStudy (retireed):

1 and 3 are multiplying

OpenStudy (retireed):

So that means 2 and 4 are already wrong, so that leaves us with .... 1) 6.71 grams, because (4.02 ⋅ 1.67) ⋅ (10^23 ⋅ 10^−24) = 6.7134 3) 0.67 grams, because (4.02 ⋅ 1.67) ⋅ (10^23 ⋅ 10^−24) = 6.7134 ⋅ 10^−1 Now it gets a bit more complicated.

OpenStudy (missmeow):

okay, what next :)

OpenStudy (retireed):

What is the difference in the answers? I'm NOT talking about subtraction. 1) 6.71 grams, because (4.02 ⋅ 1.67) ⋅ (10^23 ⋅ 10^−24) = 6.7134 3) 0.67 grams, because (4.02 ⋅ 1.67) ⋅ (10^23 ⋅ 10^−24) = 6.7134 ⋅ 10^−1 Just look at the first numbers before the grams.

OpenStudy (missmeow):

1, has a whole number, and 3, doesn't.

OpenStudy (missmeow):

at the end, of 3, there is a scientific notation

OpenStudy (retireed):

All good answers, but not the answer I am looking for. 1) 6.71 grams 3) 0.67 grams

OpenStudy (missmeow):

I don't know then

OpenStudy (retireed):

This might help you see it better .... All good answers, but not the answer I am looking for. 1) 6.710 grams 3) 0.671 grams

OpenStudy (missmeow):

1 has a decimal after the 6, whereas 3 has a decimal before the 6

OpenStudy (retireed):

Yes

OpenStudy (retireed):

Now which one is correct? 1) 6.71 grams, because (4.02 ⋅ 1.67) ⋅ (10^23 ⋅ 10^−24) = 6.7134 3) 0.67 grams, because (4.02 ⋅ 1.67) ⋅ (10^23 ⋅ 10^−24) = 6.7134 ⋅ 10^−1 Look at the scientific notation one is 10^23 and the other is 10^-24 Correct?

OpenStudy (missmeow):

3?

OpenStudy (retireed):

You didn't answer my question.

OpenStudy (retireed):

Look at the scientific notation one is 10^23 and the other is 10^-24 Correct?

OpenStudy (missmeow):

originally yes.

OpenStudy (retireed):

Yes if you multiply 10^23 and 10^-24 You get 10^-1 SO THREE is CORRECT

OpenStudy (missmeow):

Yay! I can be taught! Thank you so much!

OpenStudy (retireed):

Hang in there cat!